PROTON-TRANSFER AND HYDROGEN-BONDING IN THE GROUND AND FIRST EXCITED SINGLET STATES OF MALONALDEHYDE

\underline{A. BRACAMONTE}, A. A. ARIAS, and P. H. VACCARO , Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520-8107 USA.

As the smallest organic molecule exhibiting a symmetric hydron migration process mediated by an intramolecular hydrogen bond, malonaldehyde (HO–CH=CH–CH=O) provides a model system for probing the influence of nuclear and electronic degrees of freedom upon unimolecular dynamics. The vibrationless level of the ground electronic potential surface (X̃^{1}A_{1} ) has been the subject of numerous experimental efforts; however, proton transfer and hydrogen bonding in excited vibrational and electronic states have not been explored extensively. By exploiting polarization-selective Degenerate Four-Wave Mixing (DFWM) spectroscopy to extract information that otherwise would be hidden in the congested room temperature spectrum of bulk (gas-phase) malonaldehyde, this presentation will discuss new evidence for the dramatic changes in hydron dynamics that accompany electronic excitation into the lowest-lying singlet manifold (Ã^{1}B_{1} ). More specifically, observed Ã^{1}B_{1} tunneling splittings are consistent with a decrease in proton-transfer efficiency upon π^{∗} ← n electron promotion while excitation of O⋯O stretching motion within the X̃^{1}A_{1} potential surface is found to increase the rate of proton transfer substantially.